# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

### v2

@@ 4.2 Evaluation
Camps et al. (2024) [138] in silico ECG CV, Purkinje network Eikonal Monte Carlo Bayesian inf
@@ 4.2.1 Surface Potential Model
The prevalence of potential map estimation on the cardiac surface can be attributed to its practicality and clinical significance. It provides a powerful and accessible means of understanding cardiac electrical behavior. Electrophysiological information reconstructed on the cardiac surfaces, such as the epicardium or endocardium, is relatively easier to obtain and is invaluable for clinical insights. Most surface potential inference work primary focused on the reconstruction of epicardial (outer) potential rather than endocardial (inner) potential. This is reasonable, as epicardial potential is more accessible for measurement compared to endocardial potential. As for the forward EP model, boundary element, finite element, and Aliev-Panfilov (AP) model have been widely used.
4.2.2 Activation/ Recovery Model
Activation and recovery time maps offer valuable static isochrone images depicting the sequence of cardiac electrical depolarization and repolarization. These isochrones can be applied in the clinic in various aspects: (a) identifying sites where activation initiates, as seen in focal arrhythmias, e.g., premature ventricular contractions (PVC) origin; (b) mapping reentry circuits, e.g., in macroreentrant atrial flutter; (c) detecting abnormalities in propagation, including areas with delayed activation [190] or those exhibiting slow or abnormal conduction, e.g., scar-related ventricular tachycardia (VT) substrate; and (d) pinpointing areas with recovery abnormalities, e.g., large recovery gradients [191, 192]. To date, the majority of validation studies have concentrated on the accuracy of activation mapping, with only a few studies assessing recovery [76, 193].
Besides, one could estimate activation parameters for personalization. The estimation of EP parameters from ECG or BSP is relatively rare, as these parameters are typically more detailed and are best captured through invasive electrophysiological studies or high-resolution mapping techniques. For example, Grandits et al. [194] employed Huber regularization to estimate simultaneously distributed fibre architecture and CVs from sparse contact recordings.
@@ 4.2.3 Transmembrane Potential Model
Using the transmembrane potential (TMP) as a cardiac source model offers a highly accurate approximation of the cardiac actual electrical source [52]. Estimating volumetric myocardial TMP dynamics is clinically significant as it unveils intricate arrhythmic details and intramural arrhythmogenic substrates. However, this necessitates complex modelling and interpretation of electrical activity within the 3D cardiac structure. Reconstructions of TMP support the identification of depolarized and repolarized regions in the heart and their temporal evolution. They also delineate areas of reduced amplitude or completely absent TMP, such as ischemic, fibrotic, or border-zone areas. Nevertheless, TMPs cannot be directly measured in the clinic, necessitating indirect validation via extracellular signals generated by TMP distributions. Surface potentials on the endocardial and epicardial surfaces, as well as activation and recovery times, can be calculated from TMP distributions and then measured in experiments.
@@ 5 Clinical Application
Inverse modelling of electrocardiographic signals can help to create a personalized virtual heart, which can be potentially applied for: (a) preoperative planning for cardiac surgery, (b) risk stratification and long-term monitoring, and (c) personalized disease insight and diagnosis.
Figure 12: Virtual heart model guided preoperative planning for the personzalized catheter ablation of atrial fibrillation (AF) patients. Here, CARTO is the clinical electroanatomic navigation system. Illustrations designed referring to Boyle et al. [3].
@@
5.1 Preoperative Planning for Cardiac Surgery
Virtual heart models enable surgeons to simulate different surgical interventions and assess their feasibility and potential outcomes in a risk-free environment. For instance, utilizing personalized virtual-heart technology enables the identification of optimal infarct-related ablation targets for VT [193, 195, 196]. This can be accomplished by performing virtual multi-site delivery of electrical stimuli (pacing) from various bi-ventricular locations. Each location attempts to induce VT from a site positioned differently relative to the infarct area. The optimal ablation lesions can then be determined in each personalized virtual heart, rendering it non-inducible for VT from any pacing location. Therefore, it can terminate not only clinically manifested or induced VTs during the procedure but also all potential VTs arising from the post-infarction substrate. This includes VTs that may emerge after the initial ablation, potentially eliminating the need for repeated procedures and providing long-term freedom from VT for patients. Similarly, this technology allows for the preoperative identification of optimal ablation targets for AF, as shown in Fig. 12. It has proved invaluable in guiding ablation strategies, enabling clinicians to target specific areas responsible for atrial arrhythmias more effectively [3]. For MI management, ECG inverse inference aids in delineating the extent and location of ischemic regions within the myocardium [74, 116]. It can also be utilized to detect pulmonary vein (PV) drivers and thus assist in identifying patients with high acute success rates to PV isolation [197]. Furthermore, surgeons can virtually implant cardiac devices such as pacemakers, defibrillators, or transcatheter valves to evaluate their placement and function. Additionally, complex surgical procedures such as coronary artery bypass grafting and valve repair or replacement can be simulated to optimize surgical strategies, minimize operative risks, and improve patient outcomes.
@@ 5.2 Risk Stratification and Long-Term Monitoring
In risk stratification, ECG inverse inference techniques contribute by unravelling intricate electrophysiological patterns that signify an increased risk of adverse cardiac events. For instance, Arevalo et al. [2] created a virtual-heart arrhythmia risk predictor (VARP) for sudden cardiac death (SCD) based on a virtual heart model, as shown in Fig. 13. The heart model integrates the ventricular geometry of the patient, structural remodelling resulting from post-MI and electrical functions from sub-cellular to organ levels. Within each heart model, a virtual multi-site delivery of electrical stimuli originating from ventricular locations at varying distances to the remodelled tissue can be performed. This comprehensive evaluation allows for a thorough assessment of the patient’s heart propensity to develop ventricular arrhythmias related to MI. The VARP has been demonstrated to significantly outperform clinical metrics in predicting future arrhythmic events. The resilient and non-intrusive VARP approach holds promise in potentially averting SCD and steering clear of unnecessary implantable cardioverter-defibrillator (ICD) placements in patients with post-MI. This is particularly valuable in monitoring patients with known cardiovascular conditions, where early signs of deterioration can be identified, enabling timely intervention and personalized adjustments to the treatment plan [198, 199]. This can be achieved by combining multiple sources (fitness trackers, wearables, and ’companion’ apps for drugs, etc.) to maintain and update cardiac digital twins and their underlying models. The integration of ECG inverse inference thus emerges as a powerful tool in risk assessment and continuous monitoring, contributing to more proactive and personalized cardiovascular care.
@@ 5.3 Personalized Disease Insight and Diagnosis
Numerous studies within the pathological validation section have delved into disease mechanisms using ECG inverse inference, proposing clinically relevant parameters [200, 201, 202]. One of the earliest extensively clinically explored applications of ECG inverse inference is in identifying the specific mechanisms underlying arrhythmia onset and/ or perpetuation for ablation therapy [52]. For instance, it was utilized to delineate the distinct activation patterns during arrhythmia (e.g., focal vs. re-entrant) [203, 204]. Additionally, it has also been employed to pinpoint the drivers of this activation [205], which have demonstrated correlation with successfully ablated regions in patients with positive outcomes and are now the focus of direct targeting in clinical validation studies [206]. Besides atria, ECG inverse inference has also been applied to identify and understand arrhythmogenic substrates for the onset and maintenance of VT [201, 207] and ventricular arrhythmias [208]. It has showcased larger ventricular electrical dyssynchrony in CRT responders, improving patient selection over standard 12-lead ECG [209, 210]. ECG inverse inference techniques are especially valuable in scenarios where invasive recordings are ethically constrained, such as with infants and pregnant women, offering insights into cardiac evolution. They can also be used for in silico clinical trials (see Section 6.3), which can particularly facilitate safe methods to explore treatment effects in patients with rare diseases and thus may allow for insights not possible in the current clinical trial practice [35]. Certain diagnostic medical devices, like CardioInsight (Medtronic, USA), based on personalized models, have transitioned into industrial application and clinical usage [35]. Within the CardioInsight mapping system, the cardiac electrical activity on its surface can be reconstructed from body surface potentials employing a customized model of the heart and torso of the target patient [211]. While these retrospective studies contribute to understanding disease mechanisms and defining potential therapeutic markers, further exploration through extensive prospective studies is crucial to establish the clinical value of these insights for individual patients and their potential role in guiding diagnosis or patient selection.

### v4

@@ 3.2 Evaluation
Camps et al. (2024) [194] in silico ECG global CV, Purkinje network Eikonal Monte Carlo Bayesian inf
Zaman et al. (2021) [136] in silico; 3 VT patients ECG tissue excitability AP Bayesian active learning
Dhamala et al. (2020) [149] in silico; 3 VT patients ECG tissue excitability AP generative VAE
Li et al. (2022) [147] 1 patient ECG global CV, EAS Eikonal conditional VAE
Konukoglu et al. (2011) [139] in silico; 1 patient BSP EAS, conductivity Eikonal polynomial chaos based Bayesian inf
Dhamala et al. (2017) [195] in silico; 2 VT patients ECG tissue excitability AP spatially adaptive surrogate model
Pezzuto et al. (2022) [153] in silico ECG EAS Eikonal GP based Bayesian optimization
The validation of the inverse problem involves a diverse set of targeted cardiac sources or EP parameters, input signal, forward EP model formulations, and inverse inference methods, as presented in Table 3 and Table 4. Note that the tables only provide a selection of representative studies and are not an exhaustive list of all available studies. One can see that BSP has been widely used for reconstructing cardiac electrical activity, while ECG is commonly used for EP parameter estimation. Several studies have shown that the number of electrodes greatly affects the accuracy of cardiac source reconstruction [48, 60]. Instead of using fixed number of electrodes, Cluitmans et al. [85] proposed determining the number based on the torso size of the individual. Alternatively, Ondrusova et al. [196] proposed a two-step single dipole based inverse inference model, initially using all electrodes and then refining with a subset selected by significance via a greedy algorithm. For EP parameter estimation, the AP and (reaction-)Eikonal models are frequently used due to their balance between computational efficiency and accurate representation of cardiac activation dynamics. The Mitchell-Schaeffer (MS) model, while less commonly used, offers simplicity and is suitable for specific scenarios requiring less complexity [46, 115]. Most studies employed computer models for in silico evaluation [50, 177] and/ or performed evaluation on the basis of torso-tank experiments with isolated animal hearts, such as rabbit [197, 95], canine [88, 79, 83], and swine [104, 198, 49, 109]. Also, there exist several in-situ animal studies, where the potential of the heart and body surfaces were simultaneously recorded in animals, such as Auckland pig or Utah/ Maastricht dog data [84, 128, 173]. There are also few studies evaluated using in-vivo human subjects, which however usually employed non-simultaneous invasive recordings, such as EGMs [193] or bipolar voltages maps [186, 142]. The comparison with the published invasive measurements in animals or humans has also been utilized as an alternative evaluation [25]. As for inference target, for ECGI, there are three major cardiac source models, i.e., surface potential models (EpiP and EndoP), activation models, and transmembrane voltage models, as shown in Fig. 7. Additionally, some studies have focused on reconstructing regions of electrical conduction abnormalities [178, 105, 188, 103] and the origins of ventricular activation [65, 180, 181]. In general, reconstruction efforts have primarily focused on the ventricles, with relatively few studies addressing atrial imaging, as specifically highlighted in Table 3. This might be due to the atrial thinner walls, irregular anatomy, and more complex electrical activity [199]. For EP parameter estimation, EAS (or Purkinje network), CV, conductivity, and tissue excitability have been mainly considered.
Figure 7: Illustration of different cardiac source models in ECGI.
@@ 3.2.1 Surface Potential Model
As summarized in Table 3, surface potentials, specifically EpiP and EndoP, are the most common targets for cardiac source reconstruction. The prevalence can be attributed to its practicality and clinical significance, with EpiP being particularly favored as it is more accessible for measurement than EndoP. Furthermore, surface potentials usually require less computational power compared to full-volume models. They are generally faster to functionalize since they deal with surface data, and often rely on established methods and datasets. Nonetheless, the reconstruction of surface potentials can suffer from non-uniqueness, particularly if the input data is not sufficiently detailed.
3.2.2 Activation Time Model
ATM is commonly used to describe the propagation of electrical activity across the heart. It is a simplified model focused on the activation phase and assumes a smooth, single wavefront, which limits its applicability to more complex cardiac events, such as reentry and arrhythmias. The functionalization of ATM is relatively quick as it focuses on timing rather than detailed voltage dynamics, simplifying the modeling process. Most studies have reconstructed the ATM on the atrial and/or ventricular surfaces from BSP or ECG [179, 111, 79, 62]. Instead, some studies have reconstructed the 3D cardiac activation sequence throughout the myocardial volume, evaluated both in silico [48, 50] and in vivo, including rabbit and canine models [95, 200, 174]. For example, Schuler et al. [50] conducted an in-silico study to identify best configurations for reducing line-of-block artifacts in ATM reconstructed from EpiP or TMP. Han et al. [95] performed an in vivo validation study by comparing reconstructed activation imaging with 3D intracardiac mapping in a group of four healthy rabbits undergoing ventricular pacing from different locations, as presented in Fig. 6. The reconstruction of ATM can be more unique in terms of identifying activation sequences but might lack the details needed to reconstruct the full electrical activity accurately.
@@ 3.2.3 Transmembrane Potential Model
Using TMP as a cardiac source model provides an accurate approximation of the cardiac electrical activity, revealing critical arrhythmic details and intramural arrhythmogenic substrates [7]. However, this approach is computationally demanding due to the need for complex 3D modeling and extensive data (volumetric mesh, fibre orientation, myocardial conductivities, etc), leading to longer setup and functionalization times. Despite these challenges, TMP reconstructions are valuable for identifying depolarized and repolarized regions and detecting ischemic, fibrotic, or border-zone areas. Since TMPs cannot be directly measured clinically, validation relies on extracellular signals generated by TMP distributions. Surface potentials and ATM can be derived from TMP distributions and then measured in experiments.
3.2.4 EP Parameter
As summarized in Table 4, the parameters to be personalized are directly tied to the forward EP models. For instance, the Eikonal model typically involves the personalization of parameters such as the EAS, CV, and tissue conductivity [100, 1, 194, 2]. For example, Camps et al. [1] used dynamic time warping (DTW) and Pearson’s correlation coefficients to align simulated ECG with clinical acquired ECG, enabling the iterative updating of EASs and CVs, as shown in Fig. 4. In contrast, the primary parameter optimized in AP models is usually tissue excitability [195, 149, 136, 118]. However, these are not the only parameters that may require optimization. Depending on the specific application and model, additional parameters such as ionic channel properties, diffusion coefficients, or APD might also need to be personalized to accurately capture patient-specific cardiac behavior [2, 119]. The choice of parameters to optimize is ultimately guided by the intended clinical application and the complexity of the model, balancing between computational feasibility and the need for precision in capturing the physiological phenomena of interest.
3.3 Result
Validating the results of ECG inverse inference is highly challenging due to the lack of a unified or standardized pipeline for ECGI or EP parameter estimation. The most commonly used evaluation metric for potential and activation models is the correlation coefficient (CC) [84, 63]. For assessing the performance of pacing site, PVC origin, or cardiac excitation origin localization, the Euclidean or geodesic distance between the exact and reconstructed locations is typically employed [111, 109, 201, 202]. As mentioned earlier, validation methods, source models, and the complexity of the forward model (e.g., homogeneous vs. inhomogeneous torso models) can vary significantly, leading to significant differences in results reported across studies. Generally, most previous studies have reported a median CC of approximately 0.7 for EpiP reconstruction, and this value can reach 0.8 to 0.9 for ATM and TMP reconstructions [109, 203]. Pacing site localization errors can range from 2.0 to 20.0 mm, with a median distance of around 10 mm [109, 105, 202]. A recent study compared the pacing site localization performance of dipole and potential-based source models using the public EDGAR dataset, reporting a median distance of 25.2 mm and 13.9 mm, respectively [202]. For PVC origin localization, they reported errors of 30.2-33.0 mm and 28.9-39.2 mm for the dipole and potential-based model, respectively. The latest results for cardiac excitation origin localization reported an error of about 23.9 mm on clinical data after fine-tuning, and 44.0 mm without fine-tuning [121]. This finding highlights the challenges of generalization when training ECGI reconstruction networks supervised by simulation data. Similar conclusions were drawn by Pilia et al. [201], who reported an error of about 1.5 mm on simulated data, which increased to 32.6 mm on unseen clinical data. ECGI has also been utilized to detect infarcted regions, achieving a average Dice score of 0.4 on simulated data [142, 188]. EP parameter estimation involves determining intrinsic cardiac properties that are less directly observable, making its evaluation more complex. Furthermore, it requires sophisticated models to capture the underlying biophysical processes. Given these challenges, its validation often requires indirect methods, such as comparing simulation results against observed clinical measurement [1]. Some studies use a combination of in-silico modeling and in vivo data, employing metrics such as root mean squared error (RMSE) and Dice coefficient for evaluation [136]. These findings highlight the significant challenges in ECG inverse inference evaluation, emphasizing the need for standardized validation methods and improved model robustness.
@@ 4 Clinical Application
While ECGI and EP parameter estimation focus on different aspects, i.e., detailed electrical imaging and personalized physiological modeling, respectively, they can be effectively integrated within the broader context of CDTs. CDTs leverage personalized data to create a comprehensive virtual heart model, encompassing detailed anatomical, electrophysiological, and dynamic information. It can not only optimize diagnostic and treatment planning, as with conventional ECGI, but also support long-term risk assessment and personalized disease management. Next, we will introduce the clinical applications of ECGI and the additional advantages brought by CDTs driven by personalized EP parameter estimation. For a more comprehensive overview of clinical applications, we refer readers to existing reviews on ECGI [7, 199] and CDTs [3, 204].
4.1 Preoperative Planning for Cardiac Surgery
The ability of ECGI to non-invasively characterize cardiac activation patterns is crucial for preoperative planning, such as catheter ablation. For example, ECGI has been effectively used to guide ablation in patients with persistent AF, achieving success rates of up to 85% of patients being free from AF 12 months post-procedure [205]. However, ablation procedures are not totally free of risks, require significant resources, and are not suitable for all AF patients [206]. Supporting the potential role of ECGI in patient selection, Meo et al. [207] demonstrated that ECGI could help identify patients likely to benefit from ablation. Similarly, Rodrigo et al. [208] showed that ECGI could stratify patients by projecting the complexity of electrical patterns onto the atrial surface, recommending ablation for those with less complex substrates. Despite these advances, clinical adoption of ECGI for AF ablation remains limited due to the incomplete understanding of AF mechanisms and ablation effects. In contrast, ECGI has gained more clinical acceptance for VT ablation [209]. Netherless, in either the atrial or ventricle EP procedure setting, the cardiac imaging used in ECGI need to be better integrated with EP to optimize its usage.
Figure 8: Virtual heart model guided preoperative planning for the personzalized catheter ablation of AF patients. Here, CARTO is the clinical electroanatomic navigation system. Illustrations designed referring to Boyle et al. [5].
CDTs expand upon the capabilities of ECGI by providing a more comprehensive simulation environment for patient-specific planning. It enable the simulation of various surgical interventions, allowing surgeons to predict outcomes, optimize strategies, and assess risks in a fully controlled, risk-free setting. For instance, it has been applied to identify optimal infarct-related ablation targets for VT [210, 211]. This can be accomplished by performing virtual multi-site delivery of electrical stimuli (pacing) from various bi-ventricular locations. Each location attempts to induce VT from a site positioned differently relative to the infarct area. The optimal ablation lesions can then be determined in each personalized virtual heart, rendering it non-inducible for VT from any pacing location. Therefore, it can terminate not only clinically manifested or induced VTs during the procedure but also all potential VTs arising from the post-infarction substrate. This includes VTs that may emerge after the initial ablation, potentially eliminating the need for repeated procedures and providing long-term freedom from VT for patients. Similarly, this technology allows for the preoperative identification of optimal ablation targets for AF [5], as shown in Fig. 8. Additionally, CDTs hold the potential for the virtual implantation of cardiac devices, such as pacemakers, defibrillators, or transcatheter valves, to evaluate their placement and function [3]. Complex surgical procedures, including coronary artery bypass grafting and valve repair or replacement, can also be simulated to optimize strategies, minimize risks, and improve patient outcomes [212, 213].
@@ 4.2 Risk Stratification and Long-Term Monitoring
By imaging and quantifying substrates associated with increased risk of adverse cardiac events, ECGI has been used for risk stratification [214]. Early studies have identified key substrates for arrhythmias, such as slow discontinuous conduction pathways [215], abnormal repolarization [216], and steep repolarization gradients [217].
CDT model provides a more comprehensive, personalized approach for risk stratification. It integrates the personalized geometry of the patient, structural remodeling resulting from diseases and electrical functions from sub-cellular to organ levels. For instance, a virtual-heart arrhythmia risk predictor (VARP) has been created for sudden cardiac death (SCD) based on a virtual heart model [4], as shown in Fig. 9. Within the model, a virtual multi-site delivery of electrical stimuli originating from ventricular locations at varying distances to the remodelled tissue can be performed. It has been demonstrated to significantly outperform clinical metrics in predicting future arrhythmic events. The virtual heart model offers significant advantages over ECGI by allowing for non-invasive, personalized predictions of SCD risk, potentially preventing unnecessary implantable cardioverter-defibrillator (ICD) placements in post myocardial infarction patients. Moreover, it enables continuous monitoring of patients with known cardiovascular conditions, allowing for early detection of deterioration and timely adjustments to treatment plans [218, 219]. By integrating data from multiple sources, such as fitness trackers, wearables, and companion apps, CDTs enhance the utility of ECG inverse inference, enabling more proactive and individualized cardiovascular care.
@@ 4.3 Personalized Disease Insight and Diagnosis
Numerous studies have demonstrated the potential of ECGI for understanding disease mechanisms and identifying clinically relevant parameters [220, 221, 222]. One of the earliest and most explored clinical applications of ECGI is in the precise identification of arrhythmia mechanisms to guide ablation therapy [7]. Also, ECGI has been used to delineate the distinct activation patterns during arrhythmia (e.g., focal vs. re-entrant) [223, 224], and to identify the primary drivers of these patterns, which often correlate with successful ablation targets [225, 73]. Beyond atrial applications, it has also been effective in detecting and characterizing arrhythmogenic substrates that contribute to the onset and maintenance of VT [221, 226] and other ventricular arrhythmias [227] It has been shown to improve patient selection for CRT by identifying greater ventricular electrical dyssynchrony among responders compared to standard 12-lead ECG methods [228, 229]. The growing commercial availability of ECGI systems, such as CardioInsight (Medtronic, USA), Corify Care, Vivo (Catheter Precision), and EP Solutions SA, further highlights its potential for clinical use [31, 75].
At the same time, CDTs are transforming the landscape of in silico clinical trials, particularly for evaluating treatment effects in patients with rare diseases, where traditional trials have significant limitations [31]. By enabling highly personalized simulations that capture complex cardiac dynamics and patient-specific characteristics, CDTs enhance the generalizability and reliability of trial results while reducing time and costs [194, 230]. This technique accelerates drug development and provides deeper insights into drug effects, interactions, and mechanisms. While these studies have significantly advanced our understanding of cardiac disease mechanisms and identified potential therapeutic targets, further prospective studies are needed to validate these insights in clinical settings and explore their potential for guiding diagnosis and patient selection.
@@ 5 Discussion and Future Perspectives
Here, we provide a structured discussion on the current challenges in ECG inverse inference and propose potential solutions to address these challenges. By evaluating uncertainties, exploring computationally efficient models, and integrating multimodal data, we aim to provide a comprehensive roadmap for future research in this field.
5.1 Quantifying and Mitigating Uncertainties in ECG Inverse Inference
Despite significant advances in techniques used for ECG inverse inference, uncertainty remains inadequately quantified across many aspects of the pipeline. Uncertainty quantification (UQ) aims to quantify the range of possible outcomes and their associated probabilities due to input variability [231]. Recently, UQ has been widely employed in cardiac EP modeling, primarily for forward problems, to examine how variations in geometry and EP parameters affect the simulated ECG morphology [232, 233, 234]. Instead, several studies have focused on how input variations inversely affect the reconstructed cardiac sources or estimated EP parameters, namely inverse UQ. These is generally achieved via polynomial chaos expansion, which allows for a compact and efficient representation of the uncertainty in model outputs [235, 236, 237]. Furthermore, probabilistic approaches, which are often used for inverse UQ, naturally trace uncertainty from observed data back to the model estimation. ECG inverse problems are typically ill-posed, often lacking unique solutions, especially in complex scenarios such as activation reconstruction or model personalization incorporating CV and Purkinje networks. While uniqueness can sometimes be recovered in standard ECG imaging (the linear case), it is less likely in these more advanced models. However, clinicians typically require a single, reliable prediction for decision-making in patient care. Therefore, it is crucial to quantify and mitigate various uncertainties to enhance the robustness and clinical applicability of ECG inverse solutions. As illustrated in the ECG inverse inference pipeline presented in Fig. 1, the primary sources of uncertainty include measurement errors, anatomical modeling inaccuracies, and assumptions in computational modeling.
ECG or BSP data typically exhibit sparsity with measured values associated with a higher degree of uncertainty. For instance, Njeru et al. [238] quantified the uncertainty on the reconstructed EpiP due to the ECG measurement. Moreover, there is ambiguity in matching the BSPM measurement points with the computational mesh due to the non-synchronized BSPM and imaging acquisition [139]. The cardiac position is also susceptible to various sources of movement artifacts, including respiration and cardiac contraction, introducing geometric errors into the ECGI solutions [239]. The geometric error in the anatomical modeling (forward model) should also be considered, as it will in turn cause estimation errors in the ECG inverse inference [125, 130, 240]. The anatomical modeling is normally started with image segmentation, so several studies directly performed UQ to investigate the effects of inaccurate segmentation on the reconstructed cardiac source [235, 241]. Incorporating a prior model can mitigate uncertainties, and a combination of prior models may be preferred for estimating complex source structures [137]. Also, one can disentangle the inter-subject anatomical variation to mitigate its effect on the ECG inverse problem [242]. Besides the heart, the torso geometry accuracy can also subtly influence the inverse solution, as the EP measurement is generally performed on the body surface [243]. There exist several studies performed UQ to investigate the effects of cardiac position inside the torso and torso conductivity on the inverse inference [244, 245, 237]. To address uncertainties, advanced regularization techniques that encompass prior knowledge, noise reduction strategies, and advanced data assimilation methods show promise [54, 246].
Model assumptions are necessary simplifications or idealizations of reality that are made to facilitate computational modeling. In the ECG inverse inference, assumptions are often made about the homogeneity of tissue properties, the behavior of electrical signals within the heart, and the relationship between measured ECG signals and underlying electrical activity [235, 247, 248]. While model assumptions are essential for simplifying complex systems, they can also introduce uncertainties and errors into the modeling process. For example, assuming uniform tissue properties throughout the heart and torso may lead to inaccuracies in the predicted electrical activity, especially in regions with significant heterogeneity [243, 249]. Similarly, assuming specific boundary conditions or physiological behaviors may not fully capture the complexity of real-world scenarios, leading to discrepancies between model predictions and observed data. ECG inverse inference also depends heavily on the assumptions on the initial distribution of heart electrical potential [231]. Determining the initial potential distribution at the onset of the cardiac cycle is particularly challenging due to the difficulty in accurately pinpointing the anatomical site of the earliest excitation in the heart, often associated with the intricate Purkinje networks [117]. Therefore, sophisticated modeling strategies are required to adequately account for the influence of these networks on ECG inverse inference [250, 194]. Once uncertainties are quantified, strategies can be employed to mitigate their impact on model predictions. The process of balancing model assumptions and uncertainty is often iterative, requiring continuous refinement and validation of models against experimental or clinical data. This iterative refinement process helps researchers to build more robust and reliable models over time.